You have decided to endow Faculty of Business, Dokuz Eylul University with a scholarship. It is expected to cost ​$6,000 per year to attend the university into perpetuity. You expect to give the university the endowment in 6 years and will accumulate it by making equal annual​ (end-of-year) deposits into an account. The rate of interest is expected to be 13​% for all future time periods.

a. How large must the endowment​ be?

b. How much must you deposit at the end of each of the next 6 years to accumulate the required​ amount?


1
Expert's answer
2021-07-13T13:24:06-0400

a) endowment will be

"PV=\frac{CF}{r}"

"=\frac{6000}{0.13}"


"=46153.8462 dollars"



b)we need to accumulate $46,153.8462 in 6 years

"FV=PV(1+i)^{6}"

"=6000(1.13)^6"

"=12491.71dollars"